OpenStudy (anonymous):

please help What does the product (-5 - 2i)(3 + 7i) equal? A) -15 - 14i B) -1 - 41i C) -29 -41i D) -1 + 41i E) -29 + 41i

OpenStudy (anonymous):

Just do regular distribution. Remember that i^2=1.

OpenStudy (ivettef365):

multiply as follows: (-5)(3) = -15 (-5)(7i) = -35i (-2i)(3)= -6i (-2i)(7i) = -14i^2

OpenStudy (ivettef365):

so you have -15-35i-6i-14i^2

OpenStudy (anonymous):

thats what i got. i see what i did wrong

OpenStudy (anonymous):

Oops, I meant i^2=-1. Sorry!

OpenStudy (ivettef365):

now combine like terms so you have -15 -41i -14i^2

OpenStudy (ivettef365):

now you substitute the i^2 = -1 like Alwayshte Booworm said

OpenStudy (anonymous):

hold on afk

OpenStudy (ivettef365):

so you have -15 -41i -14(-1) -15 -41i+14

hero (hero):

Multiply in this manner: (-5 - 2i)(3 + 7i) = -5(3 + 7i) -2i(3 + 7i) = -15 -35i - 6i - 14i^2 = -15 -41i -14(-1) = 14 - 15 - 41i = -1 - 41i

OpenStudy (ivettef365):

so combine like terms and you have -1-41i

OpenStudy (anonymous):

thanks

OpenStudy (ivettef365):

yw :)